On December 20, 2020, a company pays $40,000 for an investment in equity securities with no significant influence. On December 31, 2020, the company's year-end, the stock has a market value of $37,000. The company sells the stock in 2021 for $44,000.

On its income statement, the company reports:

A. A loss of $3,000 in 2020, and a gain of $7,000 in 2021

B. No gain or loss in 2020, and a gain of $4,000 in 2021

C. A gain of $4,000 in 2020, and no gain or loss in 2021

D. No gain or loss in 2020, and a gain of $7,000 in 2021
Answer #1

OPTION A------ A loss of $3000 in 2020, and a gain of $7000 in 2021.

Since company has invested in equity securities with no significant influence, hence cost method is applied. Under Cost Method for Available for Sale securities,  investment is reflected at its fair market value at year end and consequently unrealized gain or loss is charged through income statement .

Therefore at Year end 2020

Unrealized Loss =40000 -37000 =$3000

Year 2021

Gain = 44000 - 47000 =$7000